# Scaling Laws Meet Model Architecture: Toward Inference-Efficient LLMs

### v1

@@ Appendix E Additional Inference Evaluation Results
Figure 14: Loss vs. GQA: (left) 80M model variants; (center) 145M model variants; (right) 297M model variants. Across different model sizes, the relationship between training loss and GQA varies substantially when hidden size and the mlp-to-attention ratio are fixed. The legend denotes the hidden size of each trained model.
Appendix G More Ablation Study
@@
Figure 15: Ablation Study: (left) use multiplicative calibrations without outliers; (center) use multiplicative calibrations with outliers; (right) use additive calibrations without outliers. The outlier refers to models trained with an mlp-to-attention ratio below 0.5 or above 5. We observe that outlier data points harm the scaling law fit. Moreover, while multiplicative and additive calibrations differ in formulation, their MSE and Spearman values remain nearly identical. Dots denote the data points used for fitting, while crosses indicate the test data points.
Figure 16: Joint and non-separable calibrations: (left) use multiplicative calibrations; (right) use joint and non-separable calibrations. We observe that joint and non-separable calibrations yield higher MSE and lower Spearman scores than multiplicative calibrations, indicating inferior performance. Dots denote the data points used for fitting, while crosses indicate the test data points.
Appendix H Inference FLOPs Analysis
@@
Appendix I More Large-scale Training Results
@@
Table 6: Detailed Results on Downstream Tasks for 1B Models: In this table, we show detailed results of 1B models over 9 downstream tasks.
@@
Table 7: Detailed Results on Downstream Tasks for 3B Models: In this table, we show detailed results of 3B models over 9 downstream tasks.
@@
Figure 17: Active-Experts-to-Attn on Inference Throughput: (left) 3B-A1.1B model variants; (center) 5.3B-A1.7B model variants; (right) 8.3B-A1.5B model variants. We study the effect of the Active-Experts-to-Attention ratio on inference throughput by fixing the total number of active parameters, setting GQA to 4, and using a batch size of 2048 to reduce MoE inference variance in this figure. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 1024 input and 256 output tokens.

### v2

@@ Appendix H Detailed Throughput Statistics
Figure 24: Loss vs. GQA. (left) 80M model variants; (center) 145M model variants; (right) 297M model variants. Across different model sizes, the relationship between training loss and GQA varies substantially when hidden size and the mlp-to-attention ratio are fixed. The legend denotes the hidden size of each trained model.
Appendix J More Ablation Study
@@
Figure 25: Ablation Study. (left) use multiplicative calibrations without outliers; (center) use multiplicative calibrations with outliers; (right) use additive calibrations without outliers. The outlier refers to models trained with an mlp-to-attention ratio below 0.5 or above 5. We observe that outlier data points harm the scaling law fit. Moreover, while multiplicative and additive calibrations differ in formulation, their MSE and Spearman values remain nearly identical. Dots denote the data points used for fitting, while crosses indicate the test data points.
Figure 26: Joint and non-separable calibrations. (left) use multiplicative calibrations; (right) use joint and non-separable calibrations. We observe that joint and non-separable calibrations yield higher MSE and lower Spearman scores than multiplicative calibrations, indicating inferior performance. Dots denote the data points used for fitting, while crosses indicate the test data points.
Appendix K Inference FLOPs Analysis
@@
Appendix L More Large-scale Training Results
@@
Table 9: Detailed Results on Downstream Tasks for 1B Models. In this table, we show detailed results of 1B models over 9 downstream tasks.
@@
Table 10: Detailed Results on Downstream Tasks for 3B Models. In this table, we show detailed results of 3B models over 9 downstream tasks.
@@
Figure 27: Active-Experts-to-Attn on Inference Throughput. (left) 3B-A1.1B model variants; (center) 5.3B-A1.7B model variants; (right) 8.3B-A1.5B model variants. We study the effect of the Active-Experts-to-Attention ratio on inference throughput by fixing the total number of active parameters, setting GQA to 4, and using a batch size of 2048 to reduce MoE inference variance in this figure. All evaluations are performed using the vLLM framework Kwon et al. (2023) on a single NVIDIA Ampere 40GB A100 GPU with 1024 input and 256 output tokens.